# Evaluation of Audio Compression Codecs

### v1

@@ 6 Conclusions
We evaluated several of the most common audio-compression codecs providing insights in how audio compression techniques affect sonic perceptual quality of digital music. Our tests indicate that lossy audio-compression codecs do provide better compression performance than lossless audio compression codecs. This stems mainly from the fact that lossy audio-compression algorithms sacrifice aural perceptual quality for higher compression speeds and lower final storage size. However, the Vorbis audio codec emerged as the exception to this rule. It exhibits acceptable compression efficiency when compared to its competitors while at the same time achieving perceptual quality nearly imperceptibly different compared to respective uncompressed digital audio signal.
We also evaluated RVQGAN, an AI-based audio compression scheme. While RVQGAN’s codec performance in speed and size was impressive, its perceptual quality is severely lacking. We are sure to see more advancement in this field, especially improving the perceptual quality of future AI-driven audio compression codec developments.
Our evaluation strengthens our argument that, when choosing a digital audio compression codec, focus should not only be on the audio compression performance. Users should also consider the perceptual quality of the encoded audio signal to ensure the best possible listening experience. We compliment Vorbis as an excellent lossy audio-compression codec also able to maintain high-fidelity perceptual quality.
@@ 6.1 Future Work
Expanding the database of audio codec performance measurements can determine the consistency in compression ratio, encoding speed, and decoding speed of all audio compression codec. We welcome readers to add performance measurements to our database https://github.com/samcate/fmedia-Audio-Codec-Comparison-Database.
With the recent trend of high-fidelity music streaming services like [Spotify], perceptual quality evaluation should be conducted on streaming audio. Streaming services utilize a wide range of audio-compression schemes, which can provide interesting insights.
Future research should also focus on enhancing perceptual quality in AI-driven compression methods. By integrating psycho-acoustic principles into the compression neural network, it may be possible to achieve more accurate perceptual quality, providing low-storage size high-fidelity digital music to users.

### v2

@@ 6 Conclusions
We evaluated several of the most common audio-compression codecs providing insights in how audio-compression techniques affect sonic perceptual quality of digital music. As expected, our tests indicate that lossy audio-compression codecs do provide better compression performance than lossless audio-compression codecs. This stems mainly from the fact that lossy audio-compression algorithms sacrifice aural perceptual quality for higher compression speeds and lower final storage size. However, the Vorbis audio codec emerged as the exception to this rule. It exhibits acceptable compression efficiency when compared to its competitors while at the same time achieving perceptual quality nearly imperceptibly distinguishable compared to the respective uncompressed digital-audio signal. Finally, if storage size is not a concern, lossless audio-compression codecs such as FLAC will be provider better aural perceptual quality in reproducing the original audio signal.
We also evaluated RVQGAN, a machine-learning based audio compression scheme. While RVQGAN’s codec performance in encoding speed and final storage size was impressive, its aural perceptual quality is severely lacking. We are sure to see more advancement in this field, especially improving the perceptual quality of future AI-driven audio-compression codec developments.
Our evaluation strengthens our argument that, when choosing a digital-audio compression codec, not only audio-compression performance should be a focus. Users should also consider the perceptual quality of the encoded audio signal to ensure the best possible listening experience. We compliment Vorbis as an excellent lossy audio-compression codec, which is also able to maintain high-fidelity perceptual quality. The demonstrated consistency between objective evaluation metrics, i.e. vasic PEAQ, advanced PEAQ, and the 2f model, and established MUSHRA behavior provides further evidence that objective perceptual modeling can serve as a robust foundation for audio-codec evaluation without reliance on extensive listening panels.
@@ 6.1 Future Work
Expanding our database of audio-codec performance measurements can determine the consistency in compression ratio, encoding speed, and decoding speed of all audio-compression codecs. We welcome additional performance-measurement submissions to our database https://github.com/samcate/fmedia-Audio-Codec-Comparison-Database by the community.
With the recent trend of increased use of high-fidelity music streaming services like [undefae], perceptual quality evaluation should be extended to streaming audio as well. Streaming services utilize a wide range of audio-compression schemes, which can provide interesting insights, especially when employing objective scoring.
Future research should also focus on enhancing the perceptual quality of AI-driven compression methods. By integrating psycho-acoustic principles into a compression neural network it may be possible to achieve more accurate perceptual quality in addition to providing low-storage size high-fidelity digital music to users.